A tank having a capacity of 1000 liters, initially contains 400 liters of sugar water having a concen-

tration of 0.2 Kg of sugar for each liter of water. At time zero, sugar water with a concentration of

50 gm of sugar per liter begins pumped into the tank at a rate of 2 liter per minute. Simultaneously,

a drain is opened at the bottom of the tank so that the volume of the sugar-water solution in the

tank reduces 1 liter per minute. Determine the following:



1
Expert's answer
2022-02-03T10:39:10-0500

Let s(t) = amount, in kg of sugar at time t. Then we have

"\frac{dt}{ds}"Β = (rate of salt into tank) - (rate of salt out of tank)

"= (0.05 \frac{kg}{L} · 5 \frac{L}{min}) + (0.04 \frac{kg}{Min} · 10 \frac{L}{min}) − \frac{s kg}{1000 L} · 15 \frac{L}{min}\\
= (0.25 \frac{kg}{Min}) + (0.04 \frac{kg}{Min} ) − \frac{15s kg}{1000 min} \\"

So we get the differential equation

"\frac{ds}{
dt} = 0.65 −
\frac{15s}{
1000}\\
\frac{ds}{
dt} =\frac{
130 − 3s}{
200}"

We separate s and t to get

"\frac{1}{
130 − 3}
ds = \frac{1}{
200}
dt"

Integrate

"\int \frac{1}{130-30s}ds=-\frac{1}{30}\ln \left|130-30s\right|+C\\
\frac{1}{200}t+C\\
\implies s =
\frac{130 − C_4e^{−3t/200}}{
3}"

Since we begin with pure water, we have s(0) = 0. Substituting,

"C_4 = 130"

"s =
\frac{130 − 130e^{−3t/200}}{
3}\\
s(60) =
\frac{130 − 130e^{−3t/200}}{
3} = 25.7153"

Thus, after one hour there is about 25.72 kg of sugar in the tank.